a mail order firm sends 70% of its orders by a parcel service that delivers late 5% of the time. the remaining orders are sent by a combination of a bus service and a local transportation that delivers 9% of the time. the delivery time of one system is independent of the other. if nasee uses the services of this firm, what is the probability that any two of her orders will be delivered late?
Answer #1

here P(an order is late)=P(from parcel service and late)+P(from local transportation and late)

=0.7*0.05+0.3*0.09=0.062

therefore  probability that any two of her orders will be delivered late =0.062*0.062=0.003844